# Inferring fine-grained migration patterns across the United States

### v2

@@ Methods
First, we harmonize $E^{(t)}$ with CBG populations from the 5-year ACS estimates and the 2010 Census. We scale each row of each matrix $E^{(t)}$ such that the row sums — i.e., the CBG populations — are consistent with the constraints implied by the ACS 5-year CBG populations and the 2010 Census. For example, when we take the average sum of each row across the five matrices $\{E^{(2011)},E^{(2012)},E^{(2013)},E^{(2014)},E^{(2015)}\}$, this should be equivalent to the population for the corresponding CBG in the 2011-2015 ACS. Imposing these constraints corresponds to solving a linear system for each row using non-linear least squares (to impose the constraint that populations cannot be negative).
Harmonizing with yearly state non-movers data
Next, we harmonize $E^{(t)}$ with the counts of movers and non-movers by state from 1-year ACS data. To do this, we aggregate the off-diagonal entries of columns corresponding to each state and scale them to match the count of movers by state. We also aggregate the diagonal entries of columns corresponding to each state and scale them to match the count of non-movers by state. These scalings (which treat the population remaining within each CBG as equivalent to the population of non-movers) assume that very few people move to another location within the same CBG, an assumption substantiated by previous research. [^9] Research has shown that median distance people moved in the US was about 10 to 15 miles in the 2010-19 period, which is far more than the average 2010 CBG radius (1.7 miles) and close to the 98-th percentile of radii [66].
@@
We finally harmonize $E^{(t)}$ with county populations from 1-year PEP data. We choose to match $E^{(t)}$ to PEP populations last due to their superior reliability. In addition, PEP estimates are used internally by the Census Bureau as controls for many other datasets [67], and ensuring MIGRATE estimates agree with PEP can lead to better downstream agreement in other datasets.
To match to PEP county populations, we apply an IPF-based algorithm, alternating between two steps until convergence: (1) we aggregate the entries of $E^{(t)}$ by columns corresponding to each county and scale them to match the PEP county population at year $t$ and (2) we aggregate the entries of $E^{(t)}$ by rows corresponding to each county and scale them to match the PEP county population at year $t-1$. Our algorithm runs for $3,000$ iterations, which we verify is sufficient for convergence. Specifically, at the end of 3,000 iterations, the total update to the matrix $E^{(t)}$—that is, the $L^{1}$ distance between the resulting matrices in two subsequent iterations—falls to the order of $10^{-7}$.
[…]
Supplementary information Supplementary Information is available detailing additional methods, validations, and findings.
Declarations
[…]
Data availability MIGRATE is available upon request for non-profit research use at migrate.tech.cornell.edu.
[…]
Code availability Code to reproduce our research findings is available on GitHub: https://github.com/gsagostini/MIGRATE.
@@ Appendix A Additional details on validations
Here we provide additional details on MIGRATE validations. §A.1 provides additional validation metrics that complement Figure 2; §A.2 presents additional details on our held-out data validations; §A.3 provides details on related work documenting biases of Infutor, which is the subject of our Figure 3; and §A.4 presents additional socioeconomic and demographic dimensions on which we assessed biases.
A.1 Additional validation metrics
Fig. S1: Detailed flow validations of MIGRATE. To complement the findings shown in Figure 2, we present validations with respect to area stayers (i.e., the diagonal entries of the migration matrix), area movers (i.e. the off-diagonal entries of the migration matrix), and raw in-migration (i.e. the number of people arriving at an area). We show (a-c) Pearson correlations of estimates and selected ACS flows as well as (d-f) how much MIGRATE reduces RMSE compared to Infutor.
Figure S1 reports validation statistics for subsets of the migration matrices: in particular, area stayers (i.e., the diagonal entries of the migration matrix), area movers (i.e. the off-diagonal entries of the migration matrix), and raw in-migration (i.e. the number of people arriving at an area). We report these metrics at both the state and county level; state-to-state flows were used in the estimation procedure, county-to-county flows were held out. For all datasets, MIGRATE estimates are highly correlated with Census estimates ($\rho$ above $0.957$ in all cases) and reduce RMSE relative to the Infutor data.
A.2 Validating the estimation procedure using held-out data
@@
(a) Pearson correlation between Census data and our estimates with CBG and state data held out. Pearson correlations with MIGRATE, which is estimated using all data sources, are presented for comparison.
@@
Table S1: Validation of estimates in held-out data settings. The full model validations are repeated for comparison. For in-migration rates, metrics are weighted by populations to account for disparate county and state sizes.
We verify that our estimation procedure yields highly correlated estimates with each Census data source even when it is removed from the datasets used for estimation. For example, we remove CBG-level Census populations from our estimation datasets, and verify that our resulting estimates remain highly correlated with CBG-level data. Specifically, we conduct our full battery of validations on two additional sets of migration estimates: “CBG held-out”, which is estimated without the CBG-level population data (i.e. $E^{(t)}$ was not scaled to match ACS 5-year CBG populations); and “State held-out”, which is estimated without the state-level flow data (i.e. $E^{(t)}$ was not scaled to match ACS 1-year state-to-state flows nor state mobility rates). Table S1 reports values for the Pearson correlation across all estimates and ground-truth as well as relative error reduction from Infutor. Values for the original MIGRATE estimates, reported in Figure 2 and Figure S1, are reproduced for easy comparison.
@@ A.3 Related work documenting biases in the Infutor data
As we report in Table M1, Infutor undercounts the national US population. This is consistent with prior work; a similar finding is reported by Qian and Tan [21], Bernstein et al. [19], and Diamond et al. [17] among others. Phillips [18] additionally reports that, relative to 2010 ACS data, Infutor undercounts all state populations, but not by much, and out-of-state movers. A contrasting finding, reported by Diamond et al. [4], is that Infutor might overcount the population according to the 2000 Census in San Francisco. Such a finding might be driven by specific spatiotemporal idiosyncrasies of the San Francisco data in 2000, which is not included in the subset we analyze; however, the bulk of the literature, and all the literature which analyzes the same subset of the data we do, is consistent with our finding that Infutor undercounts the data. Diamond et al. [4] also suggest that Infutor might overcount populations due to lacking information about deaths; this is consistent with our finding that Infutor overcounts the older populations in Figure 3c.
In Figure 3, we also assess whether under- or over-represents demographic subpopulations. (Prior to conducting this assessment, we scale the Infutor population to match the overall population, so any discrepancies are not caused merely by overall undercounting.) Ramiller et al. [22] similarly study representation biases of Infutor in King County, Washington, during the 2015-2019 5-year period. They find that Infutor undercounts the county population while overrepresenting White, older, home-owning, and high-income populations. Later, they regress population errors in tract-level demographics to find that Infutor records are overrepresented in areas with higher White and high-income populations among other characteristics. We confirm most of their findings in this limited spatiotemporal scale at national level. However, the authors use the NARC3 dataset—a subset of Infutor’s address records that contains detailed demographic information—and thus, beyond the spatiotemporal scope, our analyses might be substantially different.
@@ A.4 Assessment of additional demographic biases
Fig. S2: Additional analysis of socioeconomic and demographic biases at CBG population level. Variables are taken from the 5-year ACS CBG estimates.
We documented biases in the Infutor dataset geographically (Figure 3a) and on the dimensions of urbanization, age, race, household tenure, and poverty (Figure 3c). Here, we provide additional analyses of demographic bias by sex, educational attainment, and number of children per household (Figure S2). We find that Infutor displays negligible bias on the basis of sex. However, Infutor underrepresents populations with less formal education while overrepresenting populations with more formal education. Populations living in households with no children are overrepresented while those in households with children are underrepresented. This finding is consistent with the Infutor data documentation, which describes a lack of records for people under the age of 18 (therefore, children will be undercounted). MIGRATE essentially eliminates biases along all these demographic dimensions.

### v3

@@ Methods
First, we harmonize $E^{(t)}$ with CBG populations from the 5-year ACS estimates and the 2010 Census. We scale each row of each matrix $E^{(t)}$ such that the row sums — i.e., the CBG populations — are consistent with the constraints implied by the ACS 5-year CBG populations and the 2010 Census. For example, when we take the average sum of each row across the five matrices $\{E^{(2011)},E^{(2012)},E^{(2013)},E^{(2014)},E^{(2015)}\}$, this should be equivalent to the population for the corresponding CBG in the 2011-2015 ACS. We also impose a boundary constraint to match the 2010 population to that reported by the Decennial Census. Imposing these constraints corresponds to solving the non-negative least squares optimization problem
minimize $\|A\cdot\vec{x}-\vec{b}\|$
s.t. $\vec{x}\geq 0$ (2)
where
[…]
for each row (i.e. each CBG). The vector $\vec{x}$ represents the estimated population for the CBG in each year from 2009 to 2019, and we impose the constraint that $\vec{x}\geq 0$ to ensure populations are non-negative. We use these estimated CBG populations to harmonize our flow matrices, by scaling each row to match the estimated CBG population in the relevant year.
Harmonizing with yearly data on state movers and non-movers
Next, we harmonize $E^{(t)}$ with the counts of movers and non-movers by state from 1-year ACS data. To do this, we aggregate the off-diagonal entries of columns corresponding to each state and scale them to match the count of movers by state. We also aggregate the diagonal entries of columns corresponding to each state and scale them to match the count of non-movers by state. These scalings (which treat the population remaining within each CBG as equivalent to the population of non-movers) assume that very few people move to another location within the same CBG, an assumption substantiated by previous research: the median distance people moved in the US was about 10 to 15 miles in the 2010-19 period, which is far more than the average 2010 CBG radius (1.7 miles) and close to the 98th percentile of CBG radii lautz_long-distance_2023.
Mathematically, let $S^{(t)}_{k}$ be population of state $k$ in year $t$ and $R^{(t)}_{k}$ the population of state $k$ in year $t$ which did not move in the previous year. Then we multiply every diagonal entry $E^{(t)}_{xx}$ where CBG $x$ lies in state $r$ by
$\dfrac{R^{(t)}_{r}}{{\sum_{i}E_{ii}^{(t)}\cdot\mathds{1}\{\text{CBG }i\;\in\;\text{state }r\}}}$ (3)
and then multiply every off-diagonal entry $E^{(t)}_{xy}$ ($x\neq y$) where CBG $x$ lies in state $r$ and CBG $y$ lies in state $s$ by
$\dfrac{S^{(t)}_{s}-R^{(t)}_{s}}{{\sum_{i,j\;:\;i\neq j}E_{ij}^{(t)}\cdot\mathds{1}\{\text{CBG }j\;\in\;\text{state }s\}}}$ (4)
[…]
@@
Mathematically, let $F^{(t)}_{rs}$ be the number of movers between states $r$ and $s$ from year t1t-1 to year $t$. Then we multiply every entry $E^{(t)}_{xy}$ where CBG $x$ lies in state $r$ and CBG $y$ lies in state $s$ by
$\dfrac{F^{(t)}_{rs}}{\sum_{i,j}E^{(t)}_{ij}\cdot\mathds{1}\{\text{CBG }i\;\in\;\text{ state }r\}\cdot\mathds{1}\{\text{CBG }j\;\in\;\text{ state }s\}}$ (5)
whenever $F^{(t)}_{rs}\neq 0$. We scale both people who move between states ($r\neq s$) and who remain within the same state ($r=s$).
@@
Finally, we harmonize $E^{(t)}$ with county populations from 1-year PEP data. We choose to match $E^{(t)}$ to PEP populations last due to their superior reliability. In addition, PEP estimates are used internally by the Census Bureau as controls for many other datasets us_census_bureau_population_2022, and ensuring MIGRATE estimates agree with PEP can lead to better downstream agreement in other datasets.
To match to PEP county populations, we apply a classical IPF-based algorithm, alternating between two steps until convergence: (1) we aggregate the entries of $E^{(t)}$ by columns corresponding to each county and scale them to match the PEP county population at year $t$ and (2) we aggregate the entries of $E^{(t)}$ by rows corresponding to each county and scale them to match the PEP county population at year t1t-1.
More formally, let $P^{(t)}_{c}$ be the population of a given county $c$ at a given year $t$. We start with the matrix $M^{0}:=E^{(t)}$. We then perform the following updates for iterations $n=1,2,\dots N$. For odd $n$, we scale the blocks of columns of the matrix to match the county populations in year $t$. Specifically, for each CBG $y$ lying within county $q$, we multiply entries $M^{n-1}_{xy}$ by the scaling factor
$\dfrac{P^{(t)}_{q}}{\sum_{i,j}M^{n-1}_{ij}\cdot\mathds{1}\{\text{CBG }j\in\text{county }q\}}$ (6)
For even $n$, we scale blocks of rows of the matrix to match the county populations in year t1t-1. Specifically, for each CBG $x$ lying within county $p$, we multiply entries $M^{n-1}_{xy}$ by the scaling factor
$\dfrac{P^{(t-1)}_{p}}{\sum_{i,j}M^{n-1}_{ij}\cdot\mathds{1}\{\text{CBG }i\in\text{county }p\}}$ (7)
Our algorithm runs for $6,000$ iterations, which we verify is sufficient for convergence. Specifically, we track the L1 distance between the resulting matrices in two subsequent iterations.
@@ Appendix A Additional details on validations
Here we provide additional details on MIGRATE validations. We include additional details on our held-out data validations; a semi-synthetic data validation where we verify our method reconstructs a known ground-truth matrix; details on related work documenting biases of Infutor, which is the subject of our Figure 3; and additional socioeconomic and demographic dimensions on which we assessed biases, including undocumented immigrants.
A.1 Validating the estimation procedure using held-out data
@@
(a) Pearson correlation between Census data and our estimates with CBG and state data held out. Pearson correlations with MIGRATE, which is estimated using all data sources, are presented for comparison. As noted in the main text, perfect correlations between MIGRATE and all flows or stayers are partially driven by variation in area size and the predominance of non-mover counts among flows.
@@
Supplementary Table 1: Validation of estimates in held-out data settings. The full model validations are repeated for comparison. For in-migration rates, metrics are weighted by populations to account for disparate county and state sizes.
We verify that our estimation procedure yields highly correlated estimates with each Census data source even when it is removed from the datasets used for estimation. For example, we remove CBG-level Census populations from our estimation datasets, and verify that our resulting estimates remain highly correlated with CBG-level data. Specifically, we conduct our full battery of validations on two additional sets of migration estimates: “CBG held-out”, which is estimated without the CBG-level population data (i.e. $E^{(t)}$ was not scaled to match ACS 5-year CBG populations); and “State held-out”, which is estimated without the state-level flow data (i.e. $E^{(t)}$ was not scaled to match ACS 1-year state-to-state flows nor state mobility rates). Supplementary Table 1 reports values for the Pearson correlation across all estimates and ground-truth as well as relative error reduction from Infutor. Values for the original MIGRATE estimates, reported in Figure 2, are reproduced for easy comparison.
@@
A.2 Validating the harmonization procedure with semi-synthetic data
We further validate our harmonization procedure using semi-synthetic data, a standard check for statistical methods when ground-truth data is lacking balachandar2024domain, agostini2024bayesian, pierson2019inferring. Specifically, we assess how well our harmonization procedure can recover a known CBG-to-CBG flow matrix, given only a perturbed version of that matrix and the same set of marginal constraints available in our main analysis (e.g., CBG-level populations).
To generate realistic semi-synthetic data, we use the MIGRATE matrices as the ground-truth CBG-to-CBG flow matrices $M$, since these are likely more realistic than purely synthetic data generated from stylized models. These ground-truth matrices $M$ are not used as input to our harmonization method, but only as validation. From $M$, we produce the perturbed CBG-to-CBG flow matrix $E$ by adding multiplicative noise; we describe this process further below. We also compute the same marginals of $M$ that are available in our main analysis: specifically, CBG populations; state-level counts of movers and non-movers; state-to-state flows; and county populations. We use these marginals and the perturbed matrix $E$ as inputs to our harmonization method. We assess how well our method recovers the ground-truth matrix $M$ using the same metrics described in the main text (reduction in RMSE relative to using $E$ alone, and Pearson correlation).
To thoroughly test the robustness of our method, we consider two processes for generating $E$ from $M$. We now describe these processes.
Harmonizing in the presence of structured multiplicative noise
Our harmonization procedure assumes that the difference between the ground-truth matrix $M$ and the perturbed matrix $E$ can be described by a series of multiplicative scalings with a given structure (e.g., CBG-level scalings on each row) as described in the Methods Section. We thus first test how well our harmonization procedure recovers the ground-truth matrix if the perturbation does indeed match that structure. Specifically, we draw independent Log-Normal random variables that multiply (1) each row of $M$ (mirrors CBG population fitting), (2) each group of diagonal entries within the same state (mirrors state non-movers fitting), (3) each group of off-diagonal entries in columns within the same state (mirrors state movers fitting), (4) each group of entries within the same state-to-state pair (mirrors state flows fitting) and (5) each group of rows within the same county, and each group of columns within the same county (mirrors county population fitting). We draw the Log-Normal noise using standard deviations $\tau$ ranging from $0.05$ to $0.20$. We choose the upper bound on this range because it exceeds the level of variation we observe in real-world data. Specifically, across states (Figure 3a), the biases between ground-truth and Infutor populations have standard deviation 0.11; across counties, the standard deviation is 0.19, both exceeded by our upper bound of $\tau=0.20$. (We do verify, however, that results remain robust up to much larger $\tau$ values of 1.00).
We find that our harmonization method performs very strongly in this setting. For all standard deviations of noise, our inferred matrix achieves Pearson correlations above 0.99 either on the full CBG-CBG matrix or when restricting to CBG-CBG movers. The harmonized matrix also reduces RMSE (relative to the original perturbed matrix $E$) by an average of 97.7% across all years and noise levels (sd 0.06%), and by 91.9% when restricting to off-diagonal entries (sd 0.67%). These RMSE reductions remain consistently strong across the range of noise scales. The small residual error is an expected result of the optimization procedure; for example, we only fit once to CBG-level marginals to avoid overfitting, so are not guaranteed to match these marginals.
Harmonizing in the presence of real-world bias and independent noise
We next assess the performance of our harmonization procedure when the perturbation does not match the structure it assumes, but consists both of real-world bias and independent noise. Specifically, we produce the perturbed matrix $E$ from $M$ as follows:
$E_{ij}=M_{ij}\cdot\exp\left\{b\cdot\left(w_{i}+w_{j}\right)+\sigma\cdot Z_{ij}\right\}$
where $w_{i}$ is the z-scored share of the white population in CBG $i$ (capturing real-world biases, as shown in Figure 3c) and $Z_{ij}$ are i.i.d. standard normal variables. The parameters $b$ and $\sigma$ control the relative scales of the bias and the independent noise, respectively. We assess performance across a range of $b$ and $\sigma$, choosing $b$ to represent a realistic level of real-world bias based on our results in Figure 3c, and noise scales $\sigma$ up to 0.20, as above.
We would expect high levels of independent noise to make it very challenging to recover the off-diagonal entries of the ground-truth matrix. This is because the diagonal entries of the migration matrix are much bigger than the off-diagonal entries (since most people do not move); thus, the scaling factors learned by the harmonization procedure will be dominated by the diagonal entries. This is generally reasonable: if we want to estimate bias in the movers from a CBG, the bias of the entire CBG population likely serves as a good proxy. However, this logic fails under independent noise, since noise in the diagonals provides no information about noise in the off-diagonals. Importantly, the challenge of correcting for independent noise applies not just to our harmonization procedure, but to any correction procedure, since correcting such noise would require information about the ground truth for each entry.
Supplementary Fig. 1: Experimental validations of our method. Pearson correlation and reduction in RMSE achieved by our harmonization procedure on semi-synthetic data with real-world bias and independent noise. We report average metrics across all years; error bars plot standard deviation across years (from 2010-11 to 2018-19, $n=9$). Circles represent metrics computed on all entries of the flow matrix; squares represent metrics computed on movers only (off-diagonal elements). Columns assess metrics for flows at the CBG level, tract level, county level, and state level. We assess results when varying the level of bias (x-axis) and independent noise (line color).
Supplementary Figure 1 plots the Pearson correlation and reduction in RMSE for flows at every geographic scale. We compute metrics with respect to all flows and movers only (off-diagonal entries) for varying levels of noise and bias. Results are consistent with the analysis above. When considering all flows (as opposed to just off-diagonal entries), error is reduced almost completely and correlations are nearly perfect at all geographic scales (circular points). When restricting to off-diagonal entries specifically (square points) on the more granular geographic scales (tract and CBG levels), we find the largest error reductions when the scale of the independent noise is not too large relative to the bias, as expected. Pearson correlations decrease as bias and noise increase, but remain relatively high (above $0.91$ in all cases). State and county-level metrics for movers remain robust at all noise and bias scales because we have better marginal data with which to correct them (including state-to-state movers and county populations).
Collectively, these synthetic experiments show that our harmonization procedure performs well in recovering the ground-truth matrix as long as the structure of the perturbation lies within the family it is able to estimate; in contrast, purely independent noise is unsurprisingly harder to correct (both for our procedure and for any correction procedure). The family of perturbations our harmonization procedure can estimate is highly flexible (due to the number of scalings of varying shapes it incorporates) and indeed more flexible than classic iterative proportional fitting, which is widely used in similar demographic applications birkin_synthesissynthetic_1988, wong_reliability_1992, simpson_combining_2005, chang2021mobility, chang2024inferring and estimates only row and column-level scalings. The assumption that the perturbation is structured, as opposed to entirely independent for each entry, is also substantiated by our and previous findings of systematic demographic bias in Infutor data, which would not emerge under purely independent noise.
@@ A.3 Related work documenting biases in the Infutor data
As we report in Table 2, Infutor undercounts the national US population. This is consistent with prior work; a similar finding is reported by qian2021effects, bernstein2022contribution, and diamond_effects_2019 among others. phillips_measuring_2020 additionally reports that, relative to 2010 ACS data, Infutor undercounts all state populations, but not by much, and out-of-state movers. A contrasting finding, reported by diamond_effect_2020, is that Infutor might overcount the population according to the 2000 Census in San Francisco. Such a finding might be driven by specific spatiotemporal idiosyncrasies of the San Francisco data in 2000, which is not included in the subset we analyze; however, the bulk of the literature, and all the literature which analyzes the same subset of the data we do, is consistent with our finding that Infutor undercounts the data. diamond_effect_2020 also suggest that Infutor might overcount populations due to lacking information about deaths; this is consistent with our finding that Infutor overcounts the older populations in Figure 3c.
In Figure 3, we also assess whether under- or over-represents demographic subpopulations [^8] Prior to conducting this assessment, we scale the Infutor population to match the overall population, so any discrepancies are not caused merely by overall undercounting. ramiller_residential_2024 similarly study representation biases of Infutor in King County, Washington, during the 2015-2019 5-year period. They find that Infutor undercounts the county population while overrepresenting White, older, home-owning, and high-income populations. Later, they regress population errors in tract-level demographics to find that Infutor records are overrepresented in areas with higher White and high-income populations among other characteristics. We confirm most of their findings in this limited spatiotemporal scale at national level. However, the authors use the NARC3 dataset—a subset of Infutor’s address records that contains detailed demographic information—and thus, beyond the spatiotemporal scope, our analyses might be substantially different.
@@ A.4 Assessment of additional demographic biases
Supplementary Fig. 2: Additional analysis of socioeconomic and demographic biases at CBG population level. Variables are taken from the 5-year ACS CBG estimates. Bias is averaged over $n=5$ population data releases (American Community Survey 5-year estimates, 2015 through 2018) Error bars represent standard deviations across these releases.
We documented biases in the Infutor dataset geographically (Figure 3a) and along the dimensions of urbanization, age, race, household tenure, and poverty (Figure 3c). Here, we provide additional analyses of demographic bias by sex, educational attainment, number of children per household, and country of birth (Supplementary Figure 2). We find that Infutor displays negligible bias on the basis of sex. However, Infutor underrepresents populations with less formal education while overrepresenting populations with more formal education. Populations living in households with no children are overrepresented while those in households with children are underrepresented. This finding is consistent with the Infutor data documentation, which describes a lack of records for people under the age of 18 (therefore, children will be undercounted). Infutor also undercounts foreign-born residents by about 12%. MIGRATE essentially eliminates biases along all these demographic dimensions.
Biases in Representation of Immigrant Populations
Because of our finding that Infutor undercounts the foreign-born population, and previous research hypothesizing that Infutor may undercount undocumented immigrants specifically downes2023can, bernstein2022contribution, phillips_measuring_2020, we conduct further analyses investigating Infutor’s representation of undocumented immigrants. Infutor draws from many data sources — including cell phone bills, phone books, credit header files, public government records, property deeds, county property records, vehicle warranties, and data from vehicle repair and maintenance providers diamond_effects_2019, phillips2024homelessness, boar2023consumption — and some of these sources may contain data for undocumented immigrants. There were an estimated 11 million undocumented immigrants in the United States in the 2018-2022 period, about 27% of the estimated foreign-born population baker_estimates_2024.
To measure the biases of Infutor in representing the undocumented immigrant population, we investigate whether the relative population error correlates with estimates of the undocumented population. We use data on the yearly undocumented population per state provided by the Pew Research Center passel_what_2024, which is the most granular data available. As expected from the results in Figure 3a, which documents bias in Infutor data at the county level, there is geographic bias in the Infutor data at the state level as well. Infutor undercounts the population in states with larger proportions of undocumented immigrants: there is a Spearman correlation of $\rho=-0.401$ between the state-level Infutor error and the proportion of undocumented immigrants per state, averaging across the decade (Supplementary Figure 3b).
Supplementary Fig. 3: Infutor undercounts the undocumented immigrant population. We study the representation of the undocumented immigrant population by Infutor at the state level. (a) Average errors in state populations in Infutor data relative to Census data. MIGRATE estimates remove all state-level errors by construction. (b) Spearman correlation between the estimated share of undocumented immigrants in the state population (x-axis) and error in Infutor population estimates (y-axis). Infutor is more likely to undercount states with larger proportions of undocumented immigrants.
Importantly, as with other demographic biases in the Infutor dataset, harmonizing the dataset with Census data reduces biases with respect to immigrant populations as well. Census statistics do include undocumented immigrants us_census_bureau_about_2025, and so rescaling to match Census populations will mitigate Infutor’s biases. [^9] However, Census data may also be imperfect with respect to its coverage of undocumented migrants if they are, for example, less likely to respond to surveys.
In particular, MIGRATE corrects all bias at state level by construction. As shown in Supplementary Figure 2, MIGRATE also significantly reduces the bias for the overall foreign-born population.